Pat is 34 years older than his son Patrick. In 3 years, the sum of their ages will be 90. How old are they now?
Answer by robertb(5830)

Pat =x+34. In 3 years, Pat = x+37.
Patrick = x. In 3 years, Patrick = x+3.
hence x+37 + x+3 = 90.
2x + 40 = 90,
2x = 50,
x = 25, the age of Patrick,
x + 34 = 59, the age of Pat.